Solve the following equations. 
2x /3 = 18
Solve the following equations. 
2x /3 = 18

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

2x/3 = 18
2x=18×3
2x=54
x=54/2
x=27

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve it as follows:
2X/3 = 18
2 X = 18 × 3
2 X = 54
X = 27
 
Hope it helps!
Thank you ✌️